Circular No. 7324

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

V1494 AQUILAE = NOVA AQUILAE 1999 No. 2 M. Fujii, Kurashiki, Okayama, Japan, reports that spectroscopic ob-

servations (0.28-m reflector; resolution ∼ 1 nm; instrumental resolution ∼ 450 km/s) on Dec. 2.39 UT show Hα and Hβ with P-Cyg profiles; Hα emission has a width of 1200 km/s (FWHM). The spectrum is posted at

http://www1.harenet.ne.jp/~aikow/nova aql2.gif.

K. Ayani, Bisei Astronomical Observatory (BAO), writes: “I obtained low-dispersion spectra (range 400–900 nm; resolution 0.54 nm at Hα) of this nova on Dec. 2.4 UT with the BAO 1.01-m telescope (+ CCD spectrograph). Emission lines of the Balmer series, Mg II 448.1-nm (or He I 447.1-nm), and O I 777.4-nm are detected, all with P-Cyg profiles. Weak lines of Fe II 492.4- and 501.8-nm may also have P-Cyg profiles but are dominated by blueshifted absorption. By decomposing the Hα profile into a blueshifted Gaussian absorption and a Gaussian emission profile, it is determined that the FWHM and equivalent width of the emission component are ∼ 1700 km/s and 2.6 nm, respectively, and the blueshift of the absorption component with respect to the emission is ∼ 1020 km/s.”

W. Liller, Viña del Mar, Chile, reports that V1494 Aql was fainter than mag 10.5 (not visible) on an image obtained on Nov. 25.035 UT. Photometry by B. Skiff, Lowell Observatory (0.53-m reflector): Dec. 2.07, V = 5.03, b−y (Strömgren) = +0.40 (comparison star HD 188728). Additional visual magnitude estimates: Dec. 1.954, 5.5 (M. Mattei, Littleton, MA); 1.956, 5.6 (R. W. Sinnott, Chelmsford, MA); 1.96, 5.4 (G. Seronik, Belmont, MA); 2.685, 4.2 (R. J. Bouma, Groningen, The Netherlands); 2.70, 4.3 (H. Mikuž, Ljubljana, Slovenia); 2.722, 4.4 (S. Foglia, Milan, Italy); 2.732, 4.2 (K. Hornoch, Lelekovice, Czech Republic); 2.748, 4.2 (S. Baroni, Milan, Italy); 2.778, 4.0 (Baroni); 2.785, 3.9 (Baroni); 2.788, 3.7 (Baroni).

AX J1841.0–0535.8 A. Bamba and K. Koyama, Kyoto University, report: “ASCA found an

x-ray outburst in the Galactic plane on 1999 Oct. 4.2 UT, at α = 18^{h}41^{m}02^{s}, δ = −5^{o}35′.8 (equinox 2000.0; ± 20′′ at 90-percent confidence). The outburst was unusual, with a rise time of < 1000 s. The mean flux during the outburst was 10^{−10} erg cm^{−2} s^{−1} and highly variable. The spectrum was hard and heavily absorbed, typical of x-ray-binary pulsars in the Scutum arm, but we found no coherent pulsation. AX J1841.0–0535.8 coincides in position with one of the Ginga transients (Koyama et al. 1990, Nature 343, 148). We encourage follow-up observations of this source.”

1999 December 2  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green